# Learning Maximal Safe Sets Using Hypernetworks for MPC-based Local Trajectory Planning in Unknown Environments

### v2

@@ Abstract
This paper presents a novel learning-based approach for online estimation of maximal safe sets for local trajectory planning in unknown static environments. The neural representation of a set is used as the terminal set constraint for a model predictive control (MPC) local planner, resulting in improved recursive feasibility and safety. To achieve real-time performance and desired generalization properties, we employ the idea of hypernetworks. We use the Hamilton-Jacobi (HJ) reachability analysis as the source of supervision during the training process, allowing us to consider general nonlinear dynamics and arbitrary constraints. The proposed method is extensively evaluated against relevant baselines in simulations for different environments and robot dynamics. The results show a success rate increase of up to 52 % compared to the best baseline while maintaining comparable execution speed. Additionally, we deploy our proposed method, NTC-MPC, on a physical robot and demonstrate its ability to safely avoid obstacles in scenarios where the baselines fail.
@@ II RELATED WORK
One of the core problems in local motion planning is designing environment representation based on sensor data and integrating it within the planner in real-time. The representation has to provide sufficient information to safely avoid obstacles while progressing along the reference path. In practice, an Euclidian distance-based signed distance field (SDF) is a common choice due to its intuitive interpretation and efficient computation [8].
@@
The primary disadvantage of an SDF is the failure to account for the robot dynamics and actuation limits. The safe set associated with the SDF is usually an overapproximation of the maximal safe set, leading to higher collision rates [4, 6]. An alternative is to use control barrier functions (CBFs), which are based on Nagumo’s Theorem of invariance [9]. CBFs are used for different safety-critical control tasks and a discrete-time CBF was applied as a safety constraint for MPC-based trajectory planning in [10, 11]. Even though CBFs can provide desirable safety guarantees, designing a proper CBF for a general nonlinear dynamical system remains very challenging. Many authors proposed learning-based methods for CBF design [12, 13], but those approaches usually fail to characterize the safe regions well, resulting in conservative or unsafe motion.
Computing the true maximal safe regions is possible using HJ reachability analysis, which is a framework for model-based computation of reachable sets and system verification [7]. However, this framework suffers from the curse of dimensionality and is impractical for high-dimensional and time-critical systems. To improve the scalability of the method, different approaches are proposed, such as efficient initialization, system decomposition, and neural approximation of the HJ value function [14, 15, 16]. However, those approaches assume complete a priori knowledge about the environment, which is not satisfied for most real-world applications. In that regard, the work in [17] introduced parameter-conditioned reachable sets by augmenting system dynamics with a set of ”virtual” states representing environment-dependent parameters. This idea was applied for motion planning tasks in mobile robotics [18] and autonomous driving [19]. The main limitation of this method is the ability to handle only a relatively small number of environmental parameters since the dimensionality of the problem increases with the number of parameters. In contrast, our approach conditions the HJ solution based on a local observation, which can be high-dimensional, improving the existing approaches by orders of magnitude without increasing the dimensionality of the original system as in [17].
@@
Compared with other methods for safe navigation in unknown environments, the main advantage of our approach is improved real-time performance. For example, the method in [20] uses the HJ value function as the safety constraint, but it requires $\sim$600 ms to find a solution in the best case. Authors in [21] generate a neural SDF and use it as a CBF constraint for an MPC planner, requiring 150 ms for weights update and 250 ms for optimization. In contrast, depending on the horizon length and robot model our method takes between 5 and 13 ms on average, achieving significant performance improvement.
@@ III-A Problem Formulation
where $x_{r}$ is the reference state and $R$, $Q$ and $Q_{N}$ are weight matrices. We focus our attention on ${h:\mathbb{Z}\times\mathbb{R}^{n}\rightarrow\mathbb{R}}$ in 3e. The main problem addressed in this paper is how to design function $h$ in real time such that its zero-superlevel set approximates the maximal safe region.
@@ IV METHODOLOGY
From the MPC literature, it is well known that the recursive feasibility can be achieved if the terminal set is control invariant, which is true for the maximal safe set [^1] This set is also known as viability kernel. introduced above [23, 24, 25]. This motivates us to use the approximated maximal safe set as the terminal set constraint, while still using the observed SDF as the stage constraint. Nevertheless, computing maximal safe regions for general nonlinear systems under control and state constraints involves solving 9, which is intractable in real-time. Therefore, we aim for ML techniques capable of mapping directly from environment observation to the approximation of the maximal safe set in the form of a parameterized model.
@@ IV-A Model Architecture
The designed model has to satisfy two conflicting objectives: generalize well across a wide spectrum of possible observations and allow for computationally efficient trajectory optimization in real-time. Adequate generalization requires expressive models, i.e., models with a large number of trainable parameters optimized on a vast amount of data. On the other side, a computationally efficient representation requires a simple model architecture with a relatively small number of parameters. This feature is particularly important for local planners relaying on iterative numerical optimization methods such as MPC.
Our approach separates the ML model into two neural networks. The first network processes the local representation of the environment (e.g., local SDF or costmap) and outputs the parameters of the second network, which represents the constraint function in 3e. In the literature, the first network is referred to as hypernetwork, and the second network is known as main network [26]. The hypernetwork is designed to be expressive and have a large number of trainable parameters to achieve desired generalization properties. Complementary, the main network is a simple and efficient model that approximates the maximal safe set corresponding only to the current observation of the environment. During motion planning, the hypernetwork is inferred only once per time step to parametrize the main network which is then provided to the optimization algorithm as a constraint function and it is iteratively queried for its value and gradient. Moreover, the main network is naturally differentiable via backpropagation and does not require gradient approximation. As argued in [27], the overall number of trainable parameters can be orders of magnitude smaller compared to the embedding-based approach. Additionally, in V-E we analyzed the potential use of a single-network architecture to justify the benefits of the hypernetworks for real-time performance.
In our concrete implementation, we assume that the local observation is provided as an SDF. This is a common representation used in practice for mobile robots and can be efficiently computed, e.g., from an occupancy grid map. Since the input is in the form of a single-channel image, our hypernetwork consists of a deep convolutional neural network (CNN) backbone and a fully connected (FC) head. After the parameters are obtained, the main network is used as the terminal constraint, which approximates the value function in 8 and implicitly the maximal safe set. The only assumption is that the costmap is large enough so that the obstacle is visible before the robot enters the corresponding BRT. As a main network, we use a simple multi-layer perception (MLP) model. The overall architecture of our approach is presented in Fig. 2.
@@ V-A Robot Dynamics
Dynamic Unicycle: 2nd-order unicycle model used in practice to model differential drive and skid-steer types of robots. The state vector is ${x=[x_{p},\,y_{p},\,\theta,\,v,\,\omega]^{\top}}$, where $(x_{p},y_{p})$ are positional coordinates, $\theta$ is orientation, and $v$ and $\omega$ are linear and angular velocities, respectively. The control vector ${u=[a,\,\alpha]^{\top}}$ consists of linear acceleration $a$ and angular acceleration $\alpha$. The dynamic is by the following equation:
@@ V-B Model Training
The model architecture described in Section IV is implemented using PyTorch. The hypernetwork is implemented as a CNN model with single-channel input and the main network is implemented as an MLP model. The number of outputs of the hypernetwork was equal to the number of parameters of the main network. For the Dubins car model, the main network has 3 inputs, 9 hidden layers with $[32,32,32,16,16,16,8,8,8]$ neurons, and a single output, resulting in 3601 parameters in total. The first 3 layers use sinusoidal activation, while the remaining hidden layers use the SELU activation function.
The training dataset is created by first collecting a set of randomly placed local occupancy grid maps from the warehouse environment (see Fig. 4) and then computing corresponding SDFs, which are used as input to the hypernetwork. We initially collected 2.500 occupancy maps of size 6x6 m (100x100 cells with resolution 0.06 m). To augment the dataset, we perform rotations for 90, 180 and 270 degrees to every original occupancy map and to its vertically flipped version, resulting in 20.000 samples in total. Then we define the state-space grid of size 100x100x20 and compute the true value function over that grid with the numerical tool for HJ reachability in Python.
@@
During the training process, an SDF is propagated through the hypernetwork to generate parameters that are loaded into the main network, and the complete state-space grid is propagated through the main network to estimate the HJ value function over the grid. The loss function is computed based on the true and estimated value functions and its gradient is used to update the parameters of the hypernetwork (see Fig. 3). In our experiments, we use 80% of data for training and 20% for validation. We train the hypernetwork model of 4.8 million parameters for 150 epochs using Adam optimizer with the RWMSE loss defined in 12. Based on a grid search, the RWMSE hyperparameters are set to $\alpha=1000$ and $\beta=10$. For the batch size of 32, the training process takes about 7.5 hours on a single RTX3090 GPU. Fig. 5 shows the results of the model training for a sample environment observation.
@@
For the dynamic unicycle model, we perform the training process analogously using the same model architectures, except that the main network has 5 inputs. The true HJ value function is computed over a state-space grid of size 50x50x7x7x7. However, due to increased data size, the model is trained with a batch size of 16 for 100 epochs with the same optimizer and loss function for 18 h.
The hypernetwork inference time on the GPU is about $\sim$2 ms in both cases. In contrast, computing the HJ value function takes $\sim$1 s for the Dubins car and $\sim$2 s for the dynamic unicycle on the same hardware, illustrating an improvement of three orders of magnitude in execution speed.
@@ V-C Monte Carlo Simulations
for $i=0,1,\ldots,N$ and we refer to it as SDF-MPC. The second baseline called DCBF-MPC uses a safety constraint as given in [10, Example A], which resembles a discrete-time CBF. This method defines the safety constraint as
@@
We refer to our method as the neural terminal constraint MPC (NTC-MPC) local planner. All planners are implemented using the CasADi framework [28] with IPOPT as NLP solver and deployed as a ROS 2 package. The robot dynamics are discretized using the forward Euler method, and the sampling time for all MPC planners is set to 100 ms. Additionally, we use CasADi feature to autogenerate and compile C code before deployment.
To obtain a proper evaluation of our method against the baselines, we use the Monte Carlo method. For both environments and robot models, we perform 100 simulations with randomly placed obstacles of different sizes and shapes over the environments. We repeat the experiments for different lengths of the prediction horizon. In every experiment, the task of the robot is to navigate from the initial position to the goal position based only on the local observations. This scenario is common when the two successive waypoints are provided by a global planner and the robot has to navigate between them using the local planner. An experiment is considered a failure if the robot collides with an obstacle. The obtained success rates for all experiments are presented in Fig. 7. The results clearly prove an overall superior performance of our method, which is especially significant for short horizons. As expected from the theoretical discussion before, the MPC planner with neural terminal constraint representing maximal safe set avoids obstacles more successfully for short horizon lengths.
@@
Computation times reported in Fig. 8 show that our NTC-MPC method achieves real-time performance comparable to the baselines. Moreover, for horizon lengths greater than 20 steps, it is even faster than the DCBF-MPC baseline. Since we replace only the terminal constraint of the SDF-MPC with the neural network, it causes a constant offset in the computation time compared to the SDF-MPC, regardless of horizon length. On the other hand, DCBF-MPC uses constraint based on SDF increment, which is more complex than the SDF function alone, and therefore the optimization complexity grows faster when increasing the horizon length.
@@ V-D Real-World Experiments
In addition to the simulation experiments, we perform a set of collision avoidance experiments with a physical robot. For the experiments, we use a Continental delivery mobile robot which is modeled as the Dubins car model described by 13 in Section V-A. The robot uses a 3D LiDAR for local environment observations and wheel encoders for odometry. All computation is performed onboard with a mini PC Intel NUC11PHi7 (CPU: i7-1165G7 8x2.80GHz; RAM: 32GB; GPU: RTX2060 6GB).
As a representative scenario, we consider a task in which the local planner needs to avoid a static obstacle between two waypoints based only on the local observations. Like in the simulations, we conduct the experiments for different MPC methods and different lengths of the prediction horizon. The planners and the hypernetwork trained on synthetic data are directly transferred from the simulations to the real robot without any additional fine-tuning. The results of the experiments for different horizon lengths are presented in Table I, while Fig. 9 illustrates robot motion for the prediction horizon of 10 steps (1 s). Our method successfully avoids collision for all tested horizon lengths while still achieving comparable real-time performance, which is consistent with our simulations and theoretical discussion. On the other hand, the baseline MPC local planners avoid collision only if the prediction horizon is long enough.
Figure 9: Robot motion for the Dubins car model and prediction horizon of 10 time steps (1 s). Both baselines (left and middle) fail due to a short prediction horizon, while our method (right) successfully avoids collisions.
TABLE I: Results of the Collision Avoidance Experiments for the Dubins Car Model with the Physical Robot.
Results format: collision avoided [yes or not] / mean opt. time [ms]
Method $\downarrow$ |||| Horizon length $\rightarrow$ 5 10 20 30
SDF-MPC not / 1.41 not / 2.42 not / 2.74 yes / 8.21
DCBF-MPC not / 2.39 not / 2.93 yes / 6.68 yes / 8.22
NTC-MPC (ours) yes / 4.35 yes / 5.65 yes / 9.85 yes / 10.50

### v3

@@ Abstract
This paper presents a novel learning-based approach for online estimation of maximal safe sets for local trajectory planning in unknown static environments. The neural representation of a set is used as the terminal set constraint for a model predictive control (MPC) local planner, resulting in improved recursive feasibility and safety. To achieve real-time performance and desired generalization properties, we employ the idea of hypernetworks. We use the Hamilton-Jacobi (HJ) reachability analysis as the source of supervision during the training process, allowing us to consider general nonlinear dynamics and arbitrary constraints. The proposed method is extensively evaluated against relevant baselines in simulations for different environments and robot dynamics. The results show an increase in success rate of up to 52% compared to the best baseline while maintaining comparable execution speed. Additionally, we deploy our proposed method, NTC-MPC, on a physical robot and demonstrate its ability to safely avoid obstacles in scenarios where the baselines fail.
Index Terms:
Collision Avoidance, Machine Learning for Robot Control, Robot Safety
@@ II Related Work
One of the core problems in local motion planning is designing an environment representation based on sensor data and integrating it within the planner in real-time. The representation has to provide sufficient information to safely avoid obstacles while progressing along the reference path. In practice, an Euclidean distance-based signed distance function (SDF) is a common choice due to its intuitive interpretation and efficient computation [8].
@@
The primary disadvantage of an SDF is the failure to account for the robot dynamics and actuation limits. The safe set associated with the SDF is usually an overapproximation of the maximal safe set, leading to higher collision rates [4, 6]. An alternative is to use control barrier functions (CBFs), which are based on Nagumo’s Theorem of invariance [9]. CBFs are used for different safety-critical control tasks and a discrete-time CBF was applied as a safety constraint for MPC-based trajectory planning in [10]. Even though CBFs can provide desirable safety guarantees, designing a proper CBF for a general nonlinear dynamical system remains very challenging and the existing approaches usually fail to characterize the safe regions well, resulting in conservative or unsafe motion.
Computing the maximal safe set is possible using HJ reachability analysis, which is a framework for model-based computation of reachable sets and system verification [7]. This framework suffers from the curse of dimensionality and is impractical for high-dimensional systems. Improved scalability is achieved by efficient initialization [11], system decomposition [12], and neural approximation of the HJ solution [13]. However, those approaches assume complete a priori knowledge about the environment, such as a global map, to compute the HJ value function. The obtained value function would then be stored in memory and queried online during motion planning, depending on the robot’s position in the global map. If the global map is not available in advance or if the distribution of obstacles changes frequently, those methods have limited practical value.
To improve flexibility, the work in [14] introduced parameter-conditioned reachable sets by augmenting system dynamics with ”virtual” states representing environment-dependent parameters. This idea was applied for motion planning in mobile robotics [15] and autonomous driving [16]. The main limitation of this method is the ability to handle a relatively small number of environmental parameters, while our approach conditions the HJ solution based on high-dimensional observations. A different approach is used in [17] where the HJ value function is approximated numerically in a receding horizon fashion via the iterative linear-quadratic (iLQ) algorithm. Nevertheless, compared to our method, the iLQ approximation is valid only locally, the hard constraints in the system are not handled explicitly, and the resulting solution depends on initialization and system nonlinearity.
@@
In comparison with other methods for safe navigation in unknown environments, the main advantage of our approach is improved real-time performance. For example, the method in [18] uses the HJ value function as the safety constraint, but requires $\sim$600 ms to find a solution in the best case. Authors in [19] generate a neural SDF and use it as a CBF constraint for an MPC planner, requiring 150 ms for weights update and 250 ms for optimization. In contrast, depending on the horizon length and robot model, our method takes between 5 and 13 ms on average. Also, compared to MPC planners with neural constraints for collision avoidance [20] and system dynamics [21], the proposed approach uniquely employs a neural model as the MPC terminal constraint.
@@ III-A Problem Formulation
where $x_{r}$ is the reference state and $R$, $Q$ and $Q_{N}$ are weight matrices. We focus our attention on ${h:\mathbb{Z}\times\mathbb{R}^{n}\rightarrow\mathbb{R}}$ in 3e. The main problem addressed in this paper is how to design function $h$ in real-time such that its zero-superlevel set approximates the maximal safe region.
@@ IV Methodology
From the MPC literature, it is well known that the recursive feasibility can be achieved if the terminal set is control invariant, which is true for the maximal safe set [^1] This set is also known as the viability kernel. introduced above [23, 24, 25]. This motivates us to use the approximated maximal safe set as the terminal set constraint, while still using the observed SDF as the stage constraint. Nevertheless, computing maximal safe regions for general nonlinear systems under control and state constraints involves solving 9, which is intractable in real-time. Therefore, we aim for ML techniques capable of mapping directly from environment observation to the approximation of the maximal safe set in the form of a parameterized model.
@@ IV-A Model Architecture
The designed model has to satisfy two objectives: generalize well across a wide spectrum of possible observations and allow for computationally efficient trajectory optimization in real-time. Adequate generalization requires expressive models, i.e., models with a large number of trainable parameters optimized on a vast amount of data. On the other side, a computationally efficient representation requires a simple model architecture with a relatively small number of parameters. This feature is particularly important for local planners relying on iterative numerical optimization methods such as MPC.
Our approach separates the ML model into two neural networks. The first network processes the local representation of the environment (e.g., local SDF or costmap) and outputs the parameters of the second network, which represents the constraint function in 3e. In the literature, the first network is referred to as hypernetwork, and the second network is known as main network [26]. The hypernetwork is designed to be expressive and have a large number of trainable parameters to achieve desired generalization properties. Complementary, the main network is a simple and efficient model that approximates the maximal safe set corresponding only to the current observation of the environment. During motion planning, the hypernetwork is inferred only once per time step to parametrize the main network which is then provided to the optimization algorithm as a constraint function and it is iteratively queried for its value and gradient. Moreover, the main network is naturally differentiable via backpropagation and does not require gradient approximation. As shown in [27], the total number of trainable parameters can be orders of magnitude smaller compared to the embedding-based approach. Additionally, in V-E we analyzed the potential use of a single-network architecture to justify the benefits of the hypernetworks for real-time performance.
In our concrete implementation, we assume that the local observation is provided as an SDF. This is a common representation used in practice for mobile robots and can be efficiently computed, e.g., from an occupancy grid map. Since the input is in the form of a single-channel image, our hypernetwork consists of a deep convolutional neural network (CNN) backbone and a fully connected (FC) head. After the parameters are obtained, the main network is used as the terminal constraint, which approximates the value function in 8 and implicitly the maximal safe set. The only assumption is that the costmap is large enough so that the obstacle is visible before the robot enters the corresponding BRT. As a main network, we use a simple multi-layer perceptron (MLP) model. The overall architecture of our approach is presented in Fig. 2.
@@ V-A Robot Dynamics
Dynamic Unicycle: 2nd-order unicycle model used in practice to model differential drive and skid-steer types of robots. The state vector is ${x=[x_{p},\,y_{p},\,\theta,\,v,\,\omega]^{\top}}$, where $(x_{p},y_{p})$ are positional coordinates, $\theta$ is orientation, and $v$ and $\omega$ are linear and angular velocities, respectively. The control vector ${u=[a,\,\alpha]^{\top}}$ consists of linear acceleration $a$ and angular acceleration $\alpha$. The system dynamics is
@@ V-B Model Training
The model architecture described in Section IV is implemented using PyTorch. The hypernetwork is implemented as a CNN model with a single-channel input and the main network is implemented as an MLP model. The number of outputs of the hypernetwork was equal to the number of parameters of the main network. For the Dubins car model, the main network has 3 inputs, 9 hidden layers with $[32,32,32,16,16,16,8,8,8]$ neurons, and a single output, resulting in 3601 parameters in total. The first 3 layers use sinusoidal activation, while the remaining hidden layers use the SELU activation function.
The training dataset is created by first collecting a set of randomly placed local occupancy grid maps from the warehouse environment (see Fig. 4) and then computing corresponding SDFs, which are used as input to the hypernetwork. We initially collected 2.500 occupancy maps of size 6$\times$6 m (100$\times$100 cells with resolution 0.06 m). To augment the dataset, we perform rotations for 90, 180 and 270 degrees to every original occupancy map and to its vertically flipped version, resulting in 20.000 samples in total. Then we define the state-space grid of size 100$\times$100$\times$20 and compute the true value function over that grid with the hj_reachability [^3] https://github.com/StanfordASL/hj_reachability library.
@@
During the training process, an SDF is propagated through the hypernetwork to generate parameters that are loaded into the main network, and the complete state-space grid is propagated through the main network to estimate the HJ value function over the grid. The loss function is computed based on the true and estimated value functions and its gradient is used to update the parameters of the hypernetwork (see Fig. 3). In our experiments, we use 80% of data for training and 20% for validation. We train the hypernetwork model of $\sim 4.8\times 10^{6}$ parameters for 150 epochs using Adam optimizer with the RWMSE loss defined in 12. Based on a grid search, the RWMSE hyperparameters are set to $\alpha=1000$ and $\beta=10$. For the batch size of 32, the training process takes about 7.5 hours on a single Nvidia RTX3090 GPU. Fig. 5 shows the results of the model training for a sample environment observation.
@@
For the dynamic unicycle model, we perform the training process analogously using the same model architectures, except that the main network has 5 inputs. The true HJ value function is computed over a state-space grid of size 50$\times$50$\times$7$\times$7$\times$7. However, due to increased data size, the model is trained with a batch size of 16 for 100 epochs with the same optimizer and loss function for 18 h.
The hypernetwork inference time on the GPU is about $\sim$2 ms in both cases. In contrast, computing the HJ value function takes $\sim$1 s for the Dubins car and $\sim$2 s for the dynamic unicycle on the same hardware, which is an improvement of three orders of magnitude in execution speed.
@@ V-C Monte Carlo Simulations
for $i=0,1,\ldots,N$ and we refer to it as SDF-MPC. The second baseline called DCBF-MPC uses a safety constraint as given in [10, Example A], which represents a discrete-time CBF. This method defines the safety constraint as
@@
We refer to our method as the Neural Terminal Constraint MPC (NTC-MPC) local planner.
All planners are implemented using the CasADi framework [28] with IPOPT as NLP solver and deployed as a ROS 2 package. The robot dynamics are discretized using the forward Euler method with a sampling time $\delta=0.1$ s. Additionally, we use the CasADi feature to autogenerate and compile C code before deployment. In the case of NTC-MPC, the hypernetwork is implemented as a PyTorch model, while the main network is implemented using CasADi functions so it can be efficiently integrated into the computational graph during optimization. Also, the parameters of the main network are treated as time-varying parameters of the underlying NLP problem, which is solved at every MPC iteration. Once a new SDF is available, the hypernetwork is inferred to estimate those parameters, and they are directly passed to the NLP solver along with the SDF, estimated state, and the reference state.
For both environments and robot models, we perform 100 simulations with randomly placed obstacles of different sizes and shapes over the environments. The experiments are repeated for different MPC methods and different horizon lengths. Moreover, to examine the impact of the proposed RWMSE loss on the success rate, we test the NTC-MPC method both for MSE and RWMSE loss functions. The task of the robot is to navigate from the initial position to the goal position based only on the local observations. This scenario is common when the two successive waypoints are provided by a global planner and the robot has to navigate between them using the local planner. An experiment is considered a failure if the robot collides with an obstacle or gets stuck. The obtained success rates are presented in Fig. 7 and clearly illustrate better performance of the proposed method. As expected from the theoretical discussion, the NTC-MPC works well even for short prediction horizons, and the results show that the RWMSE loss function indeed has a positive impact on the success rates. Besides that, one could notice that the SDF-MPC method improves significantly when increasing the prediction horizon. However, this trend is ultimately bounded by the size of the local costmap, and increasing the MPC horizon further would not bring considerable improvement without increasing the robot’s perception field.
@@
Computation times reported in Fig. 8 show that our NTC-MPC method achieves real-time performance comparable to the baselines. Moreover, for horizon lengths greater than 20 steps, it is even faster than the DCBF-MPC baseline. Since we replace only the terminal constraint of the SDF-MPC with the neural network, it causes a constant offset in the computation time compared to the SDF-MPC, regardless of horizon length. On the other hand, DCBF-MPC uses constraints based on SDF increment, which is more complex than the SDF function alone, and therefore the optimization complexity grows faster when increasing the horizon length.
@@ V-D Real-World Experiments
In addition to the simulation experiments, we perform a set of collision avoidance experiments with a physical robot. For the experiments, we use a Continental delivery mobile robot which is modeled as the Dubins car described by 13 in Section V-A. The robot uses a 3D LiDAR for localization and local costmap computation. All computation is performed onboard with a mini PC Intel NUC11PHi7 (CPU: i7-1165G7 8x2.80GHz; RAM: 32GB; GPU: RTX2060 6GB).
We consider a task in which the local planner needs to avoid static obstacles between two waypoints based only on the local observations. We perform experiments with different MPC methods and different horizon lengths for ten different distributions of static obstacles. The planners and the hypernetwork trained on synthetic data are directly transferred from the simulations to the real robot without any additional fine-tuning. The results of the experiments are presented in Table I, while Fig. 9 shows robot motion and plots of $SDF(x_{k})$ and $\hat{V}(x_{k+N})$ for the prediction horizon of 10 steps (1 s). The plots illustrate that keeping $\hat{V}(x_{k+N})>0$, i.e., keeping the terminal state $x_{k+N}$ within the maximal safe set, prevents collision. Our method achieves a higher success rate in all settings with comparable real-time performance, which is consistent with simulations and theoretical discussion.
Figure 9: Robot motion and plots of $SDF(x_{k})$ and $\hat{V}(x_{k+N})$ for the Dubins car model and horizon length of 10 steps (1 s). Both baselines fail due to the short horizon, while our method successfully avoids collision. The plots illustrate that keeping $\hat{V}(x_{k+N})>0$ indeed prevents collision.
TABLE I: Results from Hardware Experiments with Dubins Car Model
Success Rate [%] Mean Computation Time [ms]
N [steps] $\rightarrow$ 5 10 20 30 5 10 20 30
SDF-MPC 0 0 30 70 2.65 4.55 6.26 6.54
DCBF-MPC 0 10 70 80 4.06 5.61 5.91 7.76
NTC-MPC (ours) 80 90 90 90 4.93 5.26 6.54 6.31